A manufacturing company makes two types of water​ skis, a trick ski and a slalom ski. The relevant manufacturing data are given in the table.

​Labor-Hours per Ski

Maximum​ Labor-Hours

Department

Trick Ski

Slalom Ski

Available per Day

Fabricating

55

44

140

Finishing

1

1

30

Answer parts​ (A), (B), and​ (C) below.

​(A) If the profit on a trick ski is ​$50 and the profit on a slalom ski is $30, how many of each type of ski should be manufactured each day to realize a maximum​ profit? What is the maximum​ profit? The maximum profit is ​$____. The maximum occurs when ___ trick skis and ___ slalom skis are produced.
Answer #1

Solution :- Given that, A manufacturing company makes two types of water skis, a Erick Ski and a slalom ski Let, Quantity of+HT-+ + S = 1300 - 35 Equating the constraints, From $FO. We Met at WW1$ = 1320 557 + why s = 140 + ☺ - 0 -1 T T = 1180 --1185 Maximize profit = 50X-107.27 +30X137,869. 5-6363,5+4118.07. = 124543. 2-1245-43 Maximize profit =$-1245,431 * Quantity of T